Kubin Companys relevant range of production is 15,000 to 19,000 units. When it produces and sells 17,000 units, its averageRequired: 1. If 15,000 units are produced and sold, what is the variable cost per unit produced and sold? 2. If 19,000 units
Answer #1

5) Average fixed manufacturing cost per unit = (5.6*17000)/15000 = 6.35

6) Average fixed manufacturing cost per unit = (5.6*17000)/19000 = 5.01

7) Total amount of fixed manufacturing overhead = 5.6*17000 = 95200

8) Total amount of fixed manufacturing overhead = 95200